OpenStudy (anonymous):

If your parents borrowed $150,000 at 6% interest compounded monthly to purchase a house, the length of the loan is 30 years, and the monthly payment is determined to be $899.33; How much interest did they pay over the 30 years?

OpenStudy (anonymous):

I have the amortization formula... is that good enough?

OpenStudy (dumbcow):

don't need compound interest formula here 30 years = 360 months 360*899.33 = 323,758.80 subtract original loan amount and that is total interest paid

OpenStudy (campbell_st):

find the total repaid... n = 30 x 12 = 360 r = 6/12 = 0.5% then Interest paid = 150000(1 + 0.005)^360 - 150000